OpenStudy (anonymous):

what is the domain of the sqaure root of x^2-5x+4? So far I have x=1, 4, how do I use the inequality?

OpenStudy (anonymous):

the domain will be greater than or equal to 1

OpenStudy (anonymous):

what about the 4

OpenStudy (dumbcow):

\[x^{2}-5x+4 \ge 0\] \[(x-4)(x-1) \ge 0\] Now for the product to be positive, both factors must be positive or both must be negative critical points are x=4,x=1 If x>4, then both will be positive If x<1, then both will be negative Therefore domain is, x>=4 or x<= 1 Interval notation \[x \in (-\infty,1]\cup[4,\infty)\]